Shift Summary for DAY, Dec.24, 1999

Angela Bellavance (bellavan@rice.edu)
Fri, 24 Dec 1999 15:32:02 -0600 (CST)

Shift Summary for day shift, 24/Dec/99

On Shift: S. Glazov, A. Bellavance

Hours of beam delivered at all intensities: 3.9
Hours of beam delivered at nominal intensity: 4.2
Hours of E799 data (special runs not included): 4.3
Hours of special runs (laser scans, muon runs...): 0

Accounting of lost time: TeV problems, starting new runs

========================================================================
Runs ended this shift:
Run 15281 - 5.1M events - BEAM_799
Run 15282 - ongoing - BEAM_799

Tapes brought to FCC this shift: none

========================================================================
Accesses during shift: none
========================================================================
Overview of shift:
Beam returns shortly after shift begins. Smooth running until beam
structure suddenly gets very crappy. Beam goes away and returns again
after HOPS repair. Beam goes away again for TeV quench at D4.

========================================================================
Description of Shift:

run spill time comment
------------------------------------------------------------------------
15281 231 07:22 Beam returns!!
08:14 Ed calls. Wants to be paged when run ends.
08:30 Sydney's sister calls looking for him.
332-3 09:38 TeV abort and SY energy conservation.
09:48 Beam goes away. Ch.13:"Linac RF station #1 down.
Next estimate 30 minutes."
Stop run; page Ed. He says he needs more
statistics. Page him again after the next run
ends.
10:04 Beam returns early as we init another run.
15282 1 10:07 Run starts. Two spills missed.
6-7 10:14-7 SY energy conservation.
54 11:11 Beam structure suddenly goes to hell. Call MCR.
There is "enough mass in [the MCR] to start a
black hole" and experts are looking at it.
11:30 Beam still shitty. TRD6 trips off. SashaL.
called.
71 11:41 Call MCR again and get beam sent to dump until
problem is fixed.
11:43 Ch.13:"Problem with TeV High Order Correction
element. No beam to KTeV. No est. at this time."
And we wait.
12:56 Ch.13:"Problem with TeV HOPS and RF problems. No
beam to KTeV. No est. at this time. TeV HOPS work
in progress. Est. 30 minutes."
And we wait.
13:09 MCR calls and says beam is returning.
138 13:12 Beam returns!!
145 SEM back to nominal.
172 13:57 Beam disappears again.
14:07 Ch.13:"Recovering from TeV quench at D4.
Investigating cause."
14:11 Ch.13:"...Est. to beam: 1 hour."
14:18 Notice CsI heater temp. is getting above the
warning limits.
14:43 BobT paged about CsI heater.
15:12 BobT returns page. CsI Rib temps look okay
(bot=15 and top=19). He will talk to Ron Ray about
the "Heater" temp. He will call back.
15:30 Hand off.....
========================================================================
Follow up Notes on Previous Shift

========================================================================
Items to Follow up on Next Shift

- page Ed when a decently-sized run ends / tapes fill up.
- someone should be showing up to repair the heater in the gas shed. Dave
Northacker is in charge of getting that done.
- see what BobT says about the CsI "heater" temperature.

_______
Angela Bellavance work: (630) 840-2801
bellavan@rice.edu home: (630) 548-3412
http://www.owlnet.rice.edu/~bellavan/ page: (630) 722-6861
"It has yet to be proven that intelligence has any survival value."
- Arthur C. Clarke